OpenStudy (ilovebmth1234):

Gas A effuses 0.68 times faster than gas B. What is the molar mass of gas A if the molar mass of gas B is 17 grams? A) 56 grams B) 7.82 grams C) 36.76 grams D) 0.027 grams

OpenStudy (anonymous):

:)

OpenStudy (ilovebmth1234):

:3 thanks

OpenStudy (anonymous):

anytime

OpenStudy (ilovebmth1234):

Which statement about gas effusion is true? A) lighter gases effuse faster than heavier gases B) lighter gases effuse slower then heavier gases C) effusion is not related to the molar mass of gases D) light and heavy gases effuse at the same rate

OpenStudy (ilovebmth1234):

@Missiey

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Graham's Law, the rate of effusion of a gas is inversely proportional to the square root of its molar mass. which means means that lighter gases will effuse faster than heavier ones

OpenStudy (photon336):

off the top of my head \[\frac{ r_{1} }{ r_{2} } = \sqrt{\frac{ b }{ a }}\] molar mass of gas B molar mass of gas A

OpenStudy (anonymous):

anymore